Muon piquet report week 38
Download
1 / 6

Muon Piquet Report week 38 - PowerPoint PPT Presentation


  • 116 Views
  • Uploaded on

Muon Piquet Report week 38. Oleg Maev, 26.09.2011. LHC performance. Oleg Maev, 26.09.2011. Good week for LHCb: Delivered Lumi - ~87 pb-1, Recorded Lumi - ~ 82 pb-1 And no serious problem in MUON!. Summary of calibration test of Dialog signals width. Oleg Maev, 26.09.2011.

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' Muon Piquet Report week 38' - betha


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
Muon piquet report week 38

Muon Piquet Report week 38

Oleg Maev, 26.09.2011


Lhc performance

LHC performance

Oleg Maev, 26.09.2011

Good week for LHCb: Delivered Lumi - ~87 pb-1, Recorded Lumi - ~ 82 pb-1

And no serious problem in MUON!


Summary of calibration test of dialog signals width

Summary of calibration test of Dialog signals width.

Oleg Maev, 26.09.2011

1. Width in M5R4 ~ 21.5 ns: FILL 2140 - 6,89 pb-1

(Threshold tuned a bit on C-side)

https://lblogbook.cern.ch/MUON/2747

2. Width in M5R4 ~ 18 ns: FILLs 2144 and 2147 - 7,44 pb-1

(Threshold tuned a bit on C-side)

https://lblogbook.cern.ch/MUON/2752

https://lblogbook.cern.ch/MUON/2760

3. Width in M5R4 ~ 15 ns: FILLs 2150 and 2152 - 8,02 pb-1

last run in fill 2152 with width=3 r.u. was 102324

(Threshold tuned a bit on C-side in fill 2150 only)

https://lblogbook.cern.ch/MUON/2769

https://lblogbook.cern.ch/MUON/2772

Finally we took 5 fills (22,35 pb-1).

Always we had several very noisy channels in M5CR4


Mwpc s status from hv side

MWPC’s status from HV-side

Oleg Maev, 26.09.2011

Only one new Gap affected during last week


Dq issue https lblogbook cern ch data quality 5428

DQ issue: https://lblogbook.cern.ch/Data+Quality/5428

Oleg Maev, 26.09.2011

I couldn’t find any problem in M4R1,

but M4Q3R4 was out of time during this fill

probably due to huge noise on CMB M4C1C.

All runs affected!

Finally, it seems we can flag this run as OK.

FILL 2103 - M1 removed from data taking


New low efficient region in m5q4

New low efficient region in M5Q4

Oleg Maev, 26.09.2011

Visible since 17.09.11


ad